OpenStudy (anonymous):

please answer Solve for x: negative 3 over 2, multiplied by x minus 9 equals negative 27 −24 −12 11 12

OpenStudy (anonymous):

\[-\frac{ 3 }{ 2 } \times 12 = -18 - 9 = -27\]

OpenStudy (anonymous):

THANK YOU !!!!